OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find b if a = 21 and c = 29. A) b =40 B) B = 20 C) B= 10 D) B = 400

Parth (parthkohli):

Pythagorean *whistles* You're welcome

Parth (parthkohli):

\( \color{Black}{\Rightarrow 21^2 + x^2 = 29^2 }\)

OpenStudy (anonymous):

21²= 441

OpenStudy (anonymous):

Correct?

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 = 29^2 - 21^2 }\) \( \color{Black}{\Rightarrow x^2 = 841 - 441 }\) \( \color{Black}{\Rightarrow x^2 = 400 }\) \( \color{Black}{\Rightarrow x = \sqrt{400} = ...? }\)

OpenStudy (anonymous):

So it would be 25²+400=29²

OpenStudy (anonymous):

29²=841

Parth (parthkohli):

How did you get that 25 :/

OpenStudy (anonymous):

typo.. I meant 21²

Parth (parthkohli):

Exactly

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 = 841 - 441 = 400 }\)

OpenStudy (anonymous):

Wow, thanks, that helped alot.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \sqrt{400} }\)

OpenStudy (anonymous):

its b 20cm

Parth (parthkohli):

@harshendu_pandey maybe you would like to explain your answer :)

OpenStudy (anonymous):

don't answer for them... it's open STUDy not openAnswers... :P

Parth (parthkohli):

@Ruhhr

OpenStudy (anonymous):

from pythagoras theorum c2=a2+b2 b2=c2-a2 b2=(29)2-(21)2 b2=841-441 b2=400 b=20

Parth (parthkohli):

Good, but I had already explained the same :P

OpenStudy (anonymous):

Wait I don't understand how it's 20 when I got 400

Parth (parthkohli):

What is square root of 400?

OpenStudy (anonymous):

b2 is 400... you still need to take the quare root to get b...

Parth (parthkohli):

\( \color{Black}{\Rightarrow x \times x = 400 }\) There's one number that when is multiplied to itself, gets you 400. That number is 20 :D

Parth (parthkohli):

There are actually two solutions. -20 or 20, but a side of a triangle can't be in negative :)

Parth (parthkohli):

So there is only one solution to this answer.

OpenStudy (anonymous):

Oh, alright, got it!

Parth (parthkohli):

Hah! nice!